Please help me solve this equation:
Solve using Cramer's rule.
3x+2y+2z=1
5x-y-6z=3
2x+3y+3z=4


Thank you for your time I really appreciated

Answer by TimothyLamb(4379)

3x+2y+2z=1
5x-y-6z=3
2x+3y+3z=4
---
put the system of linear equations into standard form
---
3x+2y+2z=1
5x-y-6z=3
2x+3y+3z=4
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = -1
y = 4
z = -2
---
system matrix
3 2 2
5 -1 -6
2 3 3
---
inverse of system matrix
0.6 0 -0.4
-1.08 0.2 1.12
0.68 -0.2 -0.52
---
determinant of system matrix = 25
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---